# Hamiltonian replica exchange augmented with diffusion-based generative models and importance sampling to assess biomolecular conformational basins and barriers

### v1

@@ Conclusion
In this work, we show how conformational phase space exploration in Hamiltonian replica exchange simulations could better be exploited by a stochastic generative model (DDPM) to enhance the thermodynamic and kinetic mapping of the unperturbed system. While our work is very close in spirit from previous applications of DDPMs to temperature replica exchange simulations pioneered by the Tiwary group 22, 26, we propose here two key improvements. The first one is the extension to Hamiltonian replica exchange simulations by treating the potential energy (or a fraction of it) as a random fluctuating variable in the framework of the diffusive model. This presents the distinctive advantage of enabling the study of larger systems compared to TREM as conformational exploration and exchange amongst the replicas scales much better with system size. For systems for which both TREM and HREX simulations can reliably converge, we demonstrate how DDPM can greatly enhance the determination of FES while yielding identical results for the two generalized ensemble schemes.
Because generative models remain inherently limited by the quality of the data they have been trained on, we also realized that DDPM alone could not allow for the accurate determination of significant conformational barriers separating different (sampled) conformational basins. We hereby propose a second improvement of the method, which consists in a combination of importance sampling with replica exchange and DDPM that could be iterated to gradually improve the description of the free-energy barrier(s). We apply this methodology for a complex protein system and demonstrate that a loop opening/closing pathway consistent with a previous investigation using an adaptive string method.
@@
A critical point for this second application to realistic systems is that the biased sampling requires to know, or to identify, the key slow collective variables of the system, which seems to go against the original spirit of generalized ensemble approaches that are CV-agnostic and allow to excite all the modes of the system. However, we note that this can actually be a distinct advantage. One could specifically target a known slow mode of the system while facilitating sampling along faster (but bot necessarily fast enough to be fully sampled during plain unperturbed MD simulations) degrees of freedom. In a sense, is is very similar in spirit to hybrid schemes combining generalized ensemble and specific biasing approaches2. We also note that if relevant CVs are not known a priori for a given system, other techniques, possibly AI-based24, could be used to first identify a relevant reaction coordinate and our scheme could then be applied by biasing along this CV.
We conclude by stressing that although key limitations remain (sufficient sampling by the replica exchange, identification of reaction coordinates in more complex cases if one wants to determine accurate barriers), the training of the generative model and the subsequent generation of samples comes at a negligible computational cost compared to that of the replica exchange simulations. At worse, this therefore appears as a useful and easily-usable post-processing tool that can greatly enhance the thermodynamic interpretation and exploitation of replica exchange simulation data, but also of regular MD data, as recently shown. At best, and combined with specific targeting of relevant collective variables, this enables the determination of complex conformational pathways that are not necessarily easily determined using other computational techniques.
@@ 3 Acknowledgements
The research leading to these results has received funding from the European Research Council under the European Union’s Eighth Framework Program (H2020/2014-2020)/ERC Grant Agreement No. 757111 (G.S.). This work was also supported by the "Initiative d’Excellence" program from the French State (Grant "DYNAMO", ANR-11-LABX-0011-01 to GS). ZB is funded by a PhD Fellowship from the CFM Research Foundation. The simulations presented here benefited from a local computing platform administered by G. Letessier, and was granted access to the HPC resources of TGCC under the allocation A0130811005 made by GENCI (Grand Equipement National de Calcul Intensif).
@@ 5 Data availability
The REST2-DDPM code will be made available on Github (stirnemann-group/rest2-ddpm), together with some examples. All inputs and representative simulation trajectories will be made available on a public archive upon final publication of this work.

### v2

@@ Conclusion
In this work, we show how conformational phase space exploration in Hamiltonian replica exchange simulations could better be exploited by a stochastic generative model (DDPM) to enhance the thermodynamic and kinetic mapping of the unperturbed system. While our work is very close in spirit from previous applications of DDPMs to temperature replica exchange simulations pioneered by the Tiwary group 24, 28, 58, we propose here two key improvements. The first one is the extension to Hamiltonian replica exchange (HREX, here with the solute tempering approach REST2) simulations by treating the potential energy (or a fraction of it) as a random fluctuating variable in the framework of the diffusive model. This presents the distinctive advantage of enabling the study of larger systems compared to TREM as conformational exploration and exchange amongst the replicas scales much better with system size. For systems for which both TREM and HREX simulations can reliably converge, we demonstrate how DDPM can greatly enhance the determination of FES while yielding identical results for the two generalized ensemble schemes.
Because generative models remain inherently limited by the quality of the data they have been trained on, we also realized that DDPM alone could not allow for the accurate determination of significant conformational barriers separating different (sampled) conformational basins. These states are only very transiently sampled during the conformational exploration of the replica-exchange simulations, which does not allow for successful DDPM training. We hereby propose a second improvement of the method, which consists in a combination of importance sampling with replica exchange and DDPM that could be iterated to gradually improve the description of the free-energy barrier(s). We apply this methodology for a complex protein system and demonstrate that a loop opening/closing pathway consistent with a previous investigation using an adaptive string method.
@@
A critical point for this second application to realistic systems is that the biased sampling requires to know, or to identify, the key slow collective variables of the system, which seems to go against the original spirit of generalized ensemble approaches that are CV-agnostic and allow to excite all the modes of the system. However, we note that this can actually be a distinct advantage. One could specifically target a known slow mode of the system while facilitating sampling along faster (but not necessarily fast enough to be fully sampled during plain unperturbed MD simulations) degrees of freedom. In a sense, it is very similar in spirit to hybrid schemes combining generalized ensemble and specific biasing approaches2.
Here, we focused on two systems for which we knew relevant collective variables a priori, both for DDPM training and for biasing in cases the barrier remained inaccessible. We note however that DDPM has been shown to be effective even when using dozens of CVs24, 28, 58. Therefore, the identification of suitable reaction coordinates can be approached in two complementary ways. One option is to train the DDPM on an extended set of CVs, followed by the use of a machine learning-based technique26, or a non-ML alternative, to identify the system’s slow modes and transformation pathways. Alternatively, such techniques can first be employed to identify a relevant reaction coordinate, after which our scheme could be applied by biasing along this chosen CV while accelerating sampling along all other modes of the system.
We conclude by stressing that although key limitations remain (sufficient sampling by the replica exchange, identification of reaction coordinates in more complex cases if one wants to determine accurate barriers), the training of the generative model and the subsequent generation of samples comes at a negligible computational cost compared to that of the replica exchange simulations. The overall computational gain stems from its ability to leverage information from all replicas to enhance the accuracy of the probability distribution, and consequently the free-energy surfaces, in the reference, unperturbed replica. As such, this provides the same information as much longer replica-exchange trajectories that would not mix information between the replicas during the post-processing and analysis phases.
At worse, this therefore appears as a useful and easily-usable post-processing tool that can greatly enhance the thermodynamic interpretation and exploitation of replica exchange simulation data, but also of regular MD data, as recently shown. At best, and combined with specific targeting of relevant collective variables, this enables the determination of complex conformational pathways that are not necessarily easily determined using other computational techniques.
@@ 3 Acknowledgements
The research leading to these results has received funding from the European Research Council under the European Union’s Eighth Framework Program (H2020/2014-2020)/ERC Grant Agreement No. 757111 (G.S.). This work was also supported by the "Initiative d’Excellence" program from the French State (Grant "DYNAMO", ANR-11-LABX-0011-01 to GS). ZB was funded by a PhD Fellowship from the CFM Research Foundation. The simulations presented here benefited from access to the HPC resources of TGCC under the allocation A0150811005 granted by GENCI (Grand Equipement National de Calcul Intensif).
@@ 5 Data availability
The DDPM-REST2 code is available on Github (stirnemann-group/rest2-ddpm), together with some examples. All inputs and representative simulation trajectories are available on Zenodo, doi: 10.5281/zenodo.17072659.
6 Supporting information
Additional figures (representative chignolin conformations, exploration in low-potential replicas, comparison to long unbiased MD trajectories for CLN025, KDE-based free-energy surface reconstruction for PTP1B, committor analysis).